# Gravity versus astrophysics in black hole images and photon rings: equatorial emissions and spherically symmetric spacetimes

### v1

@@ 6 Image analysis
Here, we consider an actual intensity map on the observer’s screen and we perform 100 1D intensity cuts equally spaced along the angular direction $\psi$. On each of them, with the same radial resolutions used to ray-trace the simulated primary image and $n=1$ photon ring, we compute the positions of the peaks and the widths of the profiles that are defined hereafter.
@@ 6.1 Position of the peak and width of an intensity profile
The peak along a given angular direction is the maximum of the 1D intensity profile. So, the location of the peak for a given $\psi$ is the impact parameter $b$ corresponding to the maximum intensity. Hereafter, we list the footprints on the intensity peak position of different astrophysical configurations or spacetime geometries, gradually incorporating complexities such as emission anisotropy and redshift effects.
If $\theta_{\mathrm{B}}$ is fixed to a constant value, the disc emission towards the observer is axisymmetric [^22] The local physics of the disc always remains axisymmetric.. Under this simplification, the emissivity $j_{\nu}$ of equation (21) and consequently the emitted intensity $I_{\nu}^{\mathrm{em}}$, related to $j_{\nu}$ via equation (4.2), reach their maximum at a radius $r_{\mathrm{max}}$ computed as
@@
Given the range of parameters of Table 2, $r_{\mathrm{max}}\leq r_{\mathrm{H}}$ and $j_{\nu}$ is monotonically decreasing after $r_{\mathrm{max}}$. Therefore, all our emission profiles with $\theta_{\mathrm{B}}=\mathrm{const}$ radially decrease after $r_{\mathrm{H}}$. To put it another way, if we neglect frequency shift, that is, we put $g=1$ in equation (4.2), the observed intensity profiles $I_{\nu}^{\mathrm{obs}}$ are equal to the emitted ones, via equation (4.2), and peak, in this case, at the projected radial position of the event horizon for the $n=0$ emission and at the inner boundary of the $n=1$ lensing band, that is the $n=1$ image of the equatorial horizon as explained in Section 2, for the $n=1$ photon ring. Thus, without redshift effects, our constant-$\theta_{\mathrm{B}}$ observed peak positions can be predicted with the following ingredients:
  • The metric, which defines the event horizon, the photon shell and the geodesic motion, and consequently the related image features discussed in Section 2;
  • the inclination of the source, determining the deformations due to projection effects, see Section 2 and Appendix A.
For instance, in the case of a Schwarzschild black hole, the non-redshifted observed peak positions with $\theta_{\mathrm{B}}$ constant appear to be superposed on the solid, for $n=0$, or dashed, for $n=1$, gray lines in the plots of Figure 2 and they do not vary under change of the emission profile nor of the velocity of the disc.
Still in the non-redshifted framework, if the real value of the pitch angle $\theta_{\mathrm{B}}$, between the magnetic field vector and the emission direction in the rest frame of the emitter, is included in the computation of the emissivity $j_{\nu}$ of equation (21), the prediction of $r_{\mathrm{max}}$ is less straightforward and the axisymmetry of the emission profile is broken when the disc is not observed face-on; according to equation (18.1), the quantities needed to numerically compute $\theta_{\mathrm{B}}$ are then:
  • the emitted tangent 4-vector to the photon’s geodesic that depends both on the metric and the observer’s inclination (this results from lensing effects due to the spacetime curvature);
@@
If spectral effects are taken into account, the observed intensity $I_{\nu}^{\mathrm{obs}}$ is computed according to equations (4.2) and (4.2), so that it is both geometry-reliant and astrophysics-dependant:
@@
  • the emitted or observed 4-momentum of the photon depend on the metric and on the observer’s inclination (this is again the lensing effect due to the spacetime curvature).
@@
We define the width of an intensity profile as its full width at half maximum, taking the difference between the impact parameters, above and below the peak’s one, whose associated intensity is closer to that of the maximum intensity divided by two.
The comments hereinabove regarding the dependencies of the redshift factor shaping the observed intensity $I_{\nu}^{\mathrm{obs}}$ are clearly still valid and relevant when considering its width, but contrary to the peak position which, at least for our class of synchrotron profiles and taking a constant $\theta_{\mathrm{B}}$, is invariant under change of the emission profile when spectral effects are neglected, the width is intrinsically dependent on the radiative properties of the disc.
@@ 6.2 Impact of the metric on the image features
We saw in Sections 2 and 6 that the horizon, the critical curve and the lensing bands are key elements to interpret black hole images, even though they are purely geometric, non-observable features. For instance, if we incorporate neither lensing effects in the emissivity nor frequency shift $g$, the peak positions of all our emission profiles track the apparent horizon for the $n=0$ image and the inner boundary of the first lensing band for the $n=1$ photon ring. In order to evaluate the geometrical effect on the image formation, it is then fundamental to understand how these characteristics vary when varying the spacetime metric.
We recall that the location of the event horizon under the RZ parametrization (Rezzolla & Zhidenko 2014) is only governed by the parameter $\epsilon$ and it is given by:
@@
Thus, $r_{\mathrm{H}}(\epsilon)$ is a positive decreasing function for $-1<\epsilon$ and the Schwarzschild event horizon, $r_{\mathrm{H}}=2M$, is retrieved for $\epsilon=0$, as it is shown in the first panel of Figure 5.
The radius $b_{\mathrm{c}}$ of the critical curve on the screen, or critical impact parameter, is well known for any spherically symmetric spacetime (Perlick & Tsupko 2022); for the RZ metric it reads:
@@
From equations (5), (6) we know that $N=N(r,\epsilon,a_{i})$, $i\in\mathbb{N}$, so $r_{\mathrm{c}}$ and $b_{\mathrm{c}}$ depend on these same parameters via equation (24). In particular, if we fix all the parameters to be zero except for $\epsilon$, $r_{\mathrm{c}}(\epsilon)$ and its projected radius $b_{\mathrm{c}}(\epsilon)$ decrease on the interval ]1;0.5]]-1;0.5]] - 1 ; 0.5 ], being $r_{\mathrm{c}}=3M$ and $b_{\mathrm{c}}=3\sqrt{3}M$ for $\epsilon=0$. If only $a_{1}$ varies, $r_{\mathrm{c}}(a_{1})$ and $b_{\mathrm{c}}(a_{1})$ are decreasing functions too and the Schwarzschild values are retrieved for a vanishing $a_{1}$. See the second and third panels of Figure 5 for a visual representation. We stress that the lensing bands always encompass the critical curve, so that its displacement sets in motion their inner edges.
Also the radial position $r_{\mathrm{ms}}$ of the innermost stable circular orbit only depends on $N(r,\epsilon,a_{i})$, $i\in\mathbb{N}$ via equation (4.1.1) and it plays a role in the computation of the 4-velocity profile of the emitter $\boldsymbol{u}^{\mathrm{em}}$ for a Keplerian disc, see paragraph 4.1.1, influencing then the value of its redshift factor $g$ via equation (4.2). On the treated intervals, both $r_{\mathrm{ms}}(\epsilon)$, and $r_{\mathrm{ms}}(a_{1})$, with vanishing $a_{1}$ or $\epsilon$ respectively, are decreasing functions, see last panel of Figure 5, and $r_{\mathrm{ms}}=6M$ for a Schwarzschild black hole with $\epsilon=a_{1}=0$.
To summarise, the values of $r_{\mathrm{H}}$, $b_{\mathrm{c}}$ and $r_{\mathrm{ms}}$ for a RZ non-standard black hole with positive individual parametric deviations, with $\epsilon\leq 0.5$, are smaller than in the Schwarzschild case, while they are bigger for $-1<\epsilon<0$, $-1<a_{1}<0$. This forces direct or lensed intensity’s peaks to follow the same tendencies.
Finally, we note that all the parameters of the RZ metric influence the null geodesic motion, in such a way that the impact locus on the screen of a trajectory reaching a given point around the black hole is modified by them. However, the parameters $b_{1}$ and $a_{0}$, $b_{0}$ (when PPN constraints are satisfied) engender minor modifications on the analysed ray-traced image characteristics, as it is reported in Appendix C.
@@ 6.3 Impact of the astrophysics on the observed intensity
Figure 6: Redshift profiles in a Schwarzschild spacetime for a disc with Keplerian velocity (solid lines) or radially infalling (dashed lines) seen face-on (left panel) or with inclination $i=163^{\circ}$(right panel). The $r$, $\varphi$ coordinates are those of the emission point of the photon. On the left panel, the emitted intensity with fiducial parameters and $\theta_{\mathrm{B}}=\mathrm{const}$ (dotted line) and its corresponding redshifted profiles (teal curves) are also traced.
@@
We detail here the effects of various configurations of the disc on the observed intensity $I^{\mathrm{obs}}_{\nu}$ on which the measurements of the image analysis are performed, and on the key quantity needed to compute it, that is, the redshift factor $g$, see equations (4.2) and (4.2). We recall that when $g>1$ the observed frequency appears to be blueshifted, while if $g<1$, it is redshifted.
@@
As mentioned above, at constant $\theta_{\mathrm{B}}$ all our emission profiles $I^{\mathrm{em}}_{\nu}$ are radially decreasing in the pertinent range, therefore peaking at $r_{\mathrm{inner}}=r_{\mathrm{H}}$. As for the observed profile $I^{\mathrm{obs}}_{\nu}$, given that the redshift factor $g$ always tends asymptotically to $g=1$, starting from $g=0$ at $r_{H}$, as it is recalled in Appendix B, the position of the peak of the observed intensity is moved to a further radius with respect to the one of the event horizon as it is shown in the left panel of Figure 6, reaching the screen with a corresponding larger impact parameter, see Figure 2. The observed shift is of course related to the viewing angle and the shape of the velocity profile of the disc. As for the width, if the gravitational redshift kills the profile on the left side of the peak, closer to the black hole, its right decay is mainly governed by the emission falloff.
@@
If the disc is seen face-on, that is, if $i=0^{\circ}$, the redshift factor $g$ is just a function of the radial coordinate $r$, preserving the axisymmetry of the problem, and depends on the dynamics of the disc, see left panel of Figure 6: for every radial distance $r$, $g_{\text{Keplerian}}\geq g_{\text{Radial}}$ and both redshift factors tend to 1 from below, see left panel of Figure 6. Besides, photons along a given direction $\psi$ on the screen are emitted from a fixed $\varphi$ on the disc [^23] In Kerr, a given angle $\psi$ on the observer’s screen does not correspond to a specific angle around the black hole (Johnson et al. 2020)., see Figure 3. This means that the ray-traced intensity would be the same for every direction $\psi$ and, consequently, the intensity peaks are superposed to apparent isoradial curves like the ones illustrated in Figure 2 and the measured width would be constant along all the polar directions on the screen.
@@
This azimuthal symmetry of the observed intensity is lost if the dynamical [^24] For a static disc, $g$ continues to be independent of $\varphi$, because in that case only the time components of the dot products are non-vanishing.disc is seen with some inclination and the $g(r,\varphi)$ profile of course still depends on the velocity of the disc, see right panel of Figure 6. Put differently, when some inclination is present, the well-known beaming effect, causing a boosting of the apparent luminosity of the part of the source moving towards the observer, emerges. It is visualised in Figure 3, the observer being located at $\varphi=270^{\circ}$: for the radial accretion flow velocity field, the approaching side corresponds to $\varphi\in[0^{\circ};180^{\circ}]$ (upper side of the image) with a maximum at $90^{\circ}$, whereas for a counterclockwise Keplerian velocity profile, the approaching side coincides with $\varphi\in[90^{\circ};270^{\circ}]$ (left side of the image), with a maximum at $180^{\circ}$. These variations, comparable to those of $\sin\varphi$ and $-\cos\varphi$ respectively, are coherent with the dependencies in $\varphi$ in the redshift factor when assuming a Minkowski spacetime, whose derivation is provided in Appendix B.
@@
Figure 7: Positions of the $n=0$ (left panels) and $n=1$ (right panels) peaks, observed at 345 GHz and $i=163^{\circ}$, for all the possible emission configurations with fixed $\theta_{\mathrm{B}}=90^{\circ}$ in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=0.1$ (pink) and a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.1$ (violet). Each row corresponds to a different velocity: Keplerian, radial or mixed. Critical curves are represented as dashed lines.

### v2

@@ 6 Image analysis
Here, we consider an actual intensity map on the observer’s screen. We performed 100 1D intensity cuts equally spaced along the angular direction, $\psi$. On each of them, with the same radial resolutions used to ray-trace the simulated primary image and $n=1$ photon ring, we computed the positions of the peaks and the widths of the profiles which are defined hereafter.
@@ 6.1 Position of the peak and width of an intensity profile
The peak along a given angular direction is the maximum of the 1D intensity profile. So, the location of the peak for a given $\psi$ is the impact parameter, $b$, corresponding to the maximum intensity. Hereafter, we list the footprints on the intensity peak position of different astrophysical configurations or space-time geometries, gradually incorporating complexities such as emission anisotropy and redshift effects.
If $\theta_{\mathrm{B}}$ is fixed to a constant value, the disc emission towards the observer is axisymmetric. [^22] The local physics of the disc always remains axisymmetric. Under this simplification, the emissivity, $j_{\nu}$, of equation (21) and consequently the emitted intensity, $I_{\nu}^{\mathrm{em}}$, related to $j_{\nu}$ via equation (4.2), reach their maximum at a radius, $r_{\mathrm{max}}$, computed as
@@
Given the range of parameters of Table 2, $r_{\mathrm{max}}\leq r_{\mathrm{H}}$ and $j_{\nu}$ is monotonically decreasing after $r_{\mathrm{max}}$. Therefore, all our emission profiles with $\theta_{\mathrm{B}}=\mathrm{const}$ radially decrease after $r_{\mathrm{H}}$. To put it another way, if we neglect frequency shift, that is, we put $g=1$ in equation (4.2), the observed intensity profiles, $I_{\nu}^{\mathrm{obs}}$, are equal to the emitted ones, via equation (4.2), and peak, in this case, at the projected radial position of the event horizon for the $n=0$ emission and at the inner boundary of the $n=1$ lensing band, that is the $n=1$ image of the equatorial horizon as explained in Section 2, for the $n=1$ photon ring. Thus, without redshift effects, our constant-$\theta_{\mathrm{B}}$ observed peak positions can be predicted with the following ingredients:
  • The metric, which defines the event horizon, the photon shell, and the geodesic motion, and consequently the related image features discussed in Section 2;
  • the inclination of the source, determining the deformations due to projection effects (see Section 2 and Appendix A).
For instance, in the case of a Schwarzschild black hole, the non-redshifted observed peak positions with $\theta_{\mathrm{B}}$ constant appear to be superposed on the solid, for $n=0$, or dashed, for $n=1$, grey lines in the plots of Figure 2 and they do not vary under change of the emission profile nor of the velocity of the disc.
Still in the non-redshifted framework, if the real value of the pitch angle, $\theta_{\mathrm{B}}$, between the magnetic field vector and the emission direction in the rest frame of the emitter, is included in the computation of the emissivity, $j_{\nu}$, of equation (21), the prediction of $r_{\mathrm{max}}$ is less straightforward and the axisymmetry of the emission profile is broken when the disc is not observed face-on; according to equation (18.1), the quantities needed to numerically compute $\theta_{\mathrm{B}}$ are then:
  • the emitted tangent 4-vector to the photon’s geodesic which depends both on the metric and the observer’s inclination (this results from lensing effects due to the space-time curvature);
@@
If spectral effects are taken into account, the observed intensity, $I_{\nu}^{\mathrm{obs}}$, is computed according to equations (4.2) and (4.2), so that it is both geometry-reliant and astrophysics-dependant:
@@
  • the emitted or observed 4-momentum of the photon depend on the metric and on the observer’s inclination (this is again the lensing effect due to the space-time curvature).
@@
We defined the width of an intensity profile as its full width at half maximum, taking the difference between the impact parameters, above and below the peak’s one, whose associated intensity is closer to that of the maximum intensity divided by two.
The comments hereinabove regarding the dependencies of the redshift factor shaping the observed intensity, $I_{\nu}^{\mathrm{obs}}$, are clearly still valid and relevant when considering its width, but contrary to the peak position which, at least for our class of synchrotron profiles and taking a constant $\theta_{\mathrm{B}}$, is invariant under change of the emission profile when spectral effects are neglected, the width is intrinsically dependent on the radiative properties of the disc.
@@ 6.2 Impact of the metric on the image features
We have shown in Sections 2 and 6 that the horizon, the critical curve, and the lensing bands are key elements to interpret black hole images, even though they are purely geometric, non-observable features. For instance, if we do not incorporate lensing effects in the emissivity nor frequency shift, $g$, the peak positions of all our emission profiles track the apparent horizon for the $n=0$ image and the inner boundary of the first lensing band for the $n=1$ photon ring. In order to evaluate the geometrical effect on the image formation, it is then fundamental to understand how these characteristics vary when varying the space-time metric.
We recall that the location of the event horizon under the RZ parametrisation (Rezzolla & Zhidenko 2014) is only governed by the parameter $\epsilon$ and it is given by
@@
Thus, $r_{\mathrm{H}}(\epsilon)$ is a positive decreasing function for $1<\epsilon$ and the Schwarzschild event horizon, $r_{\mathrm{H}}=2M$, is retrieved for $\epsilon=0$, as it is shown in the first panel of Figure 5.
The radius of the critical curve on the screen or critical impact parameter, $b_{\mathrm{c}}$, is well known for any spherically symmetric space-time (Perlick & Tsupko 2022); for the RZ metric it reads
@@
From equations (5), (6), we know that $N=N(r,\epsilon,a_{i})$, $i\in\mathbb{N}$, so $r_{\mathrm{c}}$ and $b_{\mathrm{c}}$ depend on these same parameters via equation (24). In particular, if we fix all the parameters to be zero except for $\epsilon$, $r_{\mathrm{c}}(\epsilon)$ and its projected radius, $b_{\mathrm{c}}(\epsilon)$, decrease on the interval ]1;0.5]]-1;0.5], being $r_{\mathrm{c}}=3M$ and $b_{\mathrm{c}}=3\sqrt{3}M$ for $\epsilon=0$. If only $a_{1}$ varies, $r_{\mathrm{c}}(a_{1})$ and $b_{\mathrm{c}}(a_{1})$ are decreasing functions too and the Schwarzschild values are retrieved for a vanishing $a_{1}$. The second and third panels of Figure 5 provide a visual representation. We stress that the lensing bands always encompass the critical curve, so that its displacement sets in motion their inner edges.
Also the radial position of the ISCO, $r_{\mathrm{ms}}$, only depends on $N(r,\epsilon,a_{i})$, $i\in\mathbb{N}$, via equation (4.1.1), and it plays a role in the computation of the 4-velocity profile of the emitter, $\boldsymbol{u}^{\mathrm{em}}$, for a Keplerian disc (see paragraph 4.1.1), influencing then the value of its redshift factor, $g$, via equation (4.2). On the treated intervals, both $r_{\mathrm{ms}}(\epsilon)$ and $r_{\mathrm{ms}}(a_{1})$, with vanishing $a_{1}$ or $\epsilon$ respectively, are decreasing functions (see last panel of Figure 5), and $r_{\mathrm{ms}}=6M$ for a Schwarzschild black hole with $\epsilon=a_{1}=0$.
To summarise, the values of $r_{\mathrm{H}}$, $b_{\mathrm{c}}$, and $r_{\mathrm{ms}}$ for a RZ non-standard black hole with positive individual parametric deviations, with $\epsilon\leq 0.5$, are smaller than in the Schwarzschild case, while they are bigger for $1<\epsilon<0$, $1<a_{1}<0$. This forces direct or lensed intensity’s peaks to follow the same tendencies.
Finally, we note that all the parameters of the RZ metric influence the null geodesic motion in such a way that the impact locus on the screen of a trajectory reaching a given point around the black hole is modified by them. However, the parameters $b_{1}$ and $a_{0}$, $b_{0}$ (when PPN constraints are satisfied) engender minor modifications on the analysed ray-traced image characteristics, as it is reported in Appendix C.
@@ 6.3 Impact of the astrophysics on the observed intensity
Figure 6: Redshift profiles in a Schwarzschild space-time for a disc with Keplerian velocity (solid lines) or radially infalling (dashed lines) seen face-on (left panel) or with inclination $i=163^{\circ}$(right panel). The $r$, $\varphi$ coordinates are those of the emission point of the photon. On the left panel, the emitted intensity with fiducial parameters and $\theta_{\mathrm{B}}=\mathrm{const}$ (dotted line) and its corresponding redshifted profiles (teal curves) are also traced.
@@
We detail here the effects of various configurations of the disc on the observed intensity, $I^{\mathrm{obs}}_{\nu}$, on which the measurements of the image analysis were performed, and on the key quantity needed to compute it, that is, the redshift factor, $g$ (see equations (4.2) and (4.2)). We recall that when $g>1$ the observed frequency appears to be blueshifted, while if $g<1$, it is redshifted.
@@
As mentioned above, at constant $\theta_{\mathrm{B}}$ all our emission profiles, $I^{\mathrm{em}}_{\nu}$, are radially decreasing in the pertinent range, therefore peaking at $r_{\mathrm{inner}}=r_{\mathrm{H}}$. As for the observed profile, $I^{\mathrm{obs}}_{\nu}$, given that the redshift factor, $g$, always tends asymptotically to $g=1$, starting from $g=0$ at $r_{H}$, as it is recalled in Appendix B, the position of the peak of the observed intensity is moved to a further radius with respect to the one of the event horizon as it is shown in the left panel of Figure 6, reaching the screen with a corresponding larger impact parameter (see Figure 2). The observed shift is of course related to the viewing angle and the shape of the velocity profile of the disc. As for the width, if the gravitational redshift kills the profile on the left side of the peak, closer to the black hole, its right decay is mainly governed by the emission falloff.
@@
If the disc is seen face-on, that is, if $i=0^{\circ}$, the redshift factor, $g$, is just a function of the radial coordinate, $r$, preserving the axisymmetry of the problem, and depends on the dynamics of the disc (see left panel of Figure 6): for every radial distance, $r$, $g_{\text{Keplerian}}\geq g_{\text{Radial}}$ and both redshift factors tend to 1 from below (see left panel of Figure 6). Besides, photons along a given direction $\psi$ on the screen are emitted from a fixed $\varphi$ on the disc [^23] In Kerr, a given angle $\psi$ on the observer’s screen does not correspond to a specific angle around the black hole (Johnson et al. 2020). (see Figure 3). This means that the ray-traced intensity would be the same for every direction $\psi$ and, consequently, the intensity peaks are superposed to apparent isoradial curves like the ones illustrated in Figure 2 and the measured width would be constant along all the polar directions on the screen.
@@
This azimuthal symmetry of the observed intensity is lost if the dynamical [^24] For a static disc, $g$ continues to be independent of $\varphi$, because in that case only the time components of the dot products are non-vanishing.disc is seen with some inclination and the $g(r,\varphi)$ profile of course still depends on the velocity of the disc (see right panel of Figure 6). Put differently, when some inclination is present, the well-known beaming effect, causing a boosting of the apparent luminosity of the part of the source moving towards the observer, emerges. It is visualised in Figure 3, the observer being located at $\varphi=270^{\circ}$: for the radial accretion flow velocity field, the approaching side corresponds to $\varphi\in[0^{\circ};180^{\circ}]$ (upper side of the image) with a maximum at $90^{\circ}$, whereas for a anticlockwise Keplerian velocity profile, the approaching side coincides with $\varphi\in[90^{\circ};270^{\circ}]$ (left side of the image), with a maximum at $180^{\circ}$. These variations, comparable to those of $\sin\varphi$ and $\cos\varphi$ respectively, are coherent with the dependencies in $\varphi$ in the redshift factor when assuming a Minkowski space-time, whose derivation is provided in Appendix B.
@@
Figure 7: Positions of the $n=0$ (left panels) and $n=1$ (right panels) peaks, observed at 345 GHz and $i=163^{\circ}$, for all the possible emission configurations with fixed $\theta_{\mathrm{B}}=90^{\circ}$ in the case of a Schwarzschild black hole (green), a RZ black hole with $a_{1}|_{\,\epsilon=0}=0.1$ (pink), and a RZ black hole with $\epsilon\,|_{\,a_{1}=0}=0.1$ (violet). Each row corresponds to a different velocity: Keplerian, radial or mixed. Critical curves are represented as dashed lines.